OpenStudy (anonymous):

An arc of length 3 feet is cut off by a central angle of pi/2 radians. Find the area of the sector formed.

OpenStudy (campbell_st):

the are length formula is \[l = r \theta\] then theta is the angle subtended... pi/2 is a quadrant so 3 = pi/2 r then r = 6/pi now find the area of the sector

OpenStudy (anonymous):

you mean cut-off as the angle is bisected?